The perimeter of a triangle is 150m and the product of differences of semi perimeter and its respective sides is 496800 m. Find the area of the triangle.

Grade:9

3 Answers

Nishant Vora IIT Patna
9 years ago
Hello Student, Please find the solution
By hero’s formula,
Area of triangle =\Delta = \sqrt{s(s-a)(s-b)(s-c)}

s= 150/2 = 75
and (s-a)(s-b)(s-c) = 496800

so \Delta = \sqrt{75*496800}
\Delta = \sqrt{75*496800}
\Delta \simeq 6104 m

Meenakshi P. Sailan
9 years ago
thank you sir..
Nishant Vora IIT Patna
9 years ago
you are welcome

_________________________________________________________________________________________________________________________________________________________________________________________________________________________________________________________________________________________________________________________________________